OpenStudy (anonymous):

How many solutions will the following system of linear equations have? –2x + 4y = –7 Y = -1/2x + 5 A. 0 B. 1 C. 2 D. infinite

OpenStudy (anonymous):

@texaschic101 @Tennistar @amistre64

OpenStudy (texaschic101):

-2x + 4y = -7 4y = 2x - 7 y = 2/4x - 7/4 y = 1/2x - 7/4 and then we have : y = -1/2x + 5 different slopes means there is 1 solution